OpenStudy (anonymous):

If ycosx=xcosy find dy/dx when x=0 and y=0

OpenStudy (anonymous):

product rule

OpenStudy (anonymous):

\[-ysin(x) + \frac{dy}{dx}\cos(x) =- x \sin(y) \frac{dy}{dx}+\cos(y)\]

OpenStudy (anonymous):

etc

OpenStudy (anonymous):

i can't understand..sorry

OpenStudy (anonymous):

To finish off where elecengineer, plug 0 in for x and y. Then solve for dy/dx.